# Adjusted Connections I: Differential Cocycles for Principal Groupoid Bundles with Connection

### v1

@@ 3.3 Adjusted local Lie-algebroid-valued connections
$\left(\mathrm{d}^{\nabla}\zeta\right)^{\alpha}_{abc}-3\zeta^{\alpha}_{ad}\mathtt{r}^{d}_{\beta}\zeta^{\beta}_{bc}=0~{}.$ (3.24c)
@@
See the introduction for an extended history about these conditions; as major references there are [KS15b, general summary of adjusted algebroid theories; but curvature condition phrased differently, not as exactness condition] and [Fis21b, the exactness condition of the curvature first appeared here].
@@
for $\mathcal{N}\in\mathsf{NQMfd}$. Here, $\mathsf{p}_{\rm curv}$ is the projection (2.47), and $\mathrm{deg}(-)$ denotes the degree of the enclosed morphism. Furthermore, the condition $R^{\rm bas}_{\nabla}=0$ ensures existence of $\mathfrak{brst}^{\rm adj}(U_{X},\mathfrak{g}[1])$ as inner homomorphisms.
@@
00 0 $=\mathrm{d}E^{a}+\mathtt{r}^{a}_{\alpha}\left(F^{\alpha}-\omega_{a\beta}^{\alpha}E^{a}A^{\beta}-\tfrac{1}{2}\zeta^{\alpha}_{ab}E^{a}E^{b}\right)-\frac{\partial\mathtt{r}^{a}_{\alpha}}{\partial m^{b}}A^{\alpha}E^{b}~{},$ (3.29)
@@ 4.1 Inner action groupoids of dg-Lie groupoids
The inner action groupoid $\mathscr{A}(\mathscr{G})$ of a dg-Lie groupoid $\mathscr{G}$ is the internal homomorphism groupoid [^32] Recall the definition of $\sigma_{\Theta}(*)$ from Definition 2.13. $\underline{\mathsf{dgGrpd}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ with objects $\underline{\mathsf{hom}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ and morphisms the natural isomorphisms between the functors corresponding to $\underline{\mathsf{hom}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ in $\mathsf{dgMfd}$.
@@
$\begin{gathered}F_{1}(x,\theta_{1},\theta_{3})=F_{1}(x,\theta_{1},\theta_{2})F_{1}(x,\theta_{2},\theta_{3})~{},\\ \mathsf{t}(F_{1}(x,\theta_{1},\theta_{2}))=F_{0}(x,\theta_{1})~{},~{}~{}~{}\mbox{and}~{}~{}~{}\mathsf{s}(F_{1}(x,\theta_{1},\theta_{2}))=F_{0}(x,\theta_{2})~{}.\end{gathered}$ (4.1b)
and are hence of the form
@@
$\begin{gathered}\mathsf{t}(m,\nu;g,\gamma)=(m,\nu)~{},~{}~{}~{}\mathsf{s}(m,\nu;g,\gamma)=(m,\nu)\mathbin{\blacktriangleleft}(g,\gamma)~{},\\ (m,\nu;g_{1},\gamma_{1})\circ\bigl{(}(m,\nu)\mathbin{\blacktriangleleft}(g_{1},\gamma_{1});g_{2},\gamma_{2}\bigr{)}=\bigl{(}m,\nu;(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})\bigr{)}\\ \mathsf{e}(m,\nu)=(m,\nu;\mathsf{e}_{m},0)~{},~{}~{}~{}(m,\nu;g,\gamma)^{-1}=\left((m,\nu)\mathbin{\blacktriangleleft}(g,\gamma);(g,\gamma)^{-1}\right)\end{gathered}$ (4.9b)
@@
$(m,\nu)\mathbin{\blacktriangleleft}(g,\gamma)=\mathsf{p}_{s}\Big{(}\mathrm{Ad}\left(g^{-1};m,\nu\right)+\vartheta^{\rm tot}(g,\gamma)\Big{)}~{},$ (4.12)
where $\mathrm{Ad}$ is the adjoint action of Definition 2.9, and $\mathsf{p}_{s}$ is the projection $\mathsf{s}^{*}\mathfrak{g}\rightarrow\mathfrak{g}$. Note that both summands depend on a choice of connection, this dependence drops out in the sum.
@@ 4.2 Adjustment groupoids of Lie groupoids
Definition 4.7 (Pre-adjustment groupoid).
Consider a Lie groupoid $\mathscr{G}$ with Lie algebroid $\mathfrak{g}[1]=\mathsf{Lie}(\mathscr{G})[1]$. Then a pre-adjustment groupoid of $\mathscr{G}$ is a dg-Lie groupoid $\mathscr{A}^{\mathsf{W}}=(\mathscr{A}^{\mathsf{W}}_{1}\rightrightarrows\mathrm{T}[1]\mathfrak{g}[1])$ together with a functor $\Psi:\mathscr{A}(\mathscr{G})\rightarrow\mathscr{A}^{\mathsf{W}}$ such that we have the following pullback diagram [^34] Here and in the following, $(\mathrm{id},0)$ denotes the embedding of a (graded) manifold into its (grade-shifted) tangent bundle as zero section.:
@@
Note that in the above definition, we could have equivalently demanded a pullback diagram involving the source maps. In the description of principal $\mathscr{G}$-connections, the functor $\Psi$ will ensure that flat connections are correctly included in our description of general connections, and the pullback diagram will ensure that there are not too many gauge transformations.
The above definition is insufficient in two points. First of all, we need to identify the generators of $\mathrm{T}[1]\mathfrak{g}[1]$ mapped to curvatures, hence an adjustment on $\mathrm{T}[1]\mathfrak{g}[1]$ in the sense of Definition 3.10. Second, we need to ensure that the finite gauge transformations given by the globular structure in the pre-adjustment groupoid $\mathscr{A}^{\mathsf{W}}$ linearize to the infinitesimal gauge transformations implied by the local adjustment structure in the BRST complex $\mathfrak{brst}^{\rm adj}(-,\mathfrak{g}[1])$. This leads to the following definition.
Definition 4.8 (Adjustment groupoid).
Consider a Lie groupoid $\mathscr{G}$ with Lie algebroid $\mathfrak{g}=\mathsf{Lie}(\mathscr{G})$. An adjustment groupoid of $\mathscr{G}$ is a pre-adjustment groupoid $\mathscr{A}^{\mathsf{W}}$, together with an adjustment of the Lie algebroid of objects $\mathrm{T}[1]\mathfrak{g}[1]$ in the sense of Definition 3.10, such that the two presheaves on $\mathsf{NQMfd}$ [^35] See (3.26) for the definition of $\underline{\mathsf{hom}}^{\rm red}(\mathcal{N},\mathrm{T}[1]\mathfrak{g}[1])$.
$\mathcal{N}\mapsto\underline{\mathsf{hom}}^{\rm red}_{\nabla}(\mathcal{N},\mathrm{T}[1]\mathfrak{g}[1])~{}~{}~{}\mbox{and}~{}~{}~{}\mathcal{N}\mapsto\mathsf{Lie}\left(\mathsf{dgGrpd}\left(\mathcal{N}\rightrightarrows\mathcal{N},\mathscr{A}^{\mathsf{W}}\right)\right)$ (4.15)
are isomorphic.
@@
Remark 4.9.
Strictly speaking, the condition that turns a pre-adjustment groupoid into an adjustment is stronger than necessary. To ensure that the global gauge transformations integrate the infinitesimal ones, it would be sufficient to consider $\mathcal{N}$ of the form $\mathrm{T}[1]Y$ for $Y$ a disjoint union of contractible manifolds. However, we want to stress the fact that an adjustment is an algebraic datum on the groupoid itself that precedes the definition of a connection. In this sense, the definition we give is more natural.
4.3 Construction of adjustment groupoids
In the following, let $\mathscr{G}=(\mathsf{G}\rightrightarrows M)$ be a general Lie groupoid and $\mathfrak{g}[1]$ its Lie algebroid, regarded as a grade-shifted vector bundle. Furthermore, let $\mathcal{H}$ be a Cartan connection on $\mathscr{G}$. The adjustment groupoid for a general Lie groupoid $\mathscr{G}$ we now construct is the action groupoid of a groupoid action of the tangent groupoid $\mathrm{T}[1]\mathscr{G}$ on the tangent prolongation $\mathrm{T}[1]\mathfrak{g}[1]$ in the sense of Definition 2.23.
We are going to define such a dg-action of $\mathrm{T}[1]\mathscr{G}$ on $\mathrm{T}[1]\mathfrak{g}[1]$ along a dg-map $\psi\colon\mathrm{T}[1]\mathfrak{g}[1]\rightarrow\mathrm{T}[1]M$, and the relevant map here is the unique map arising from the lift of [^36] As before, $\mathsf{p}_{M}$ denotes a projection onto the base $M$ of a fiber bundle $E\rightarrow M$. $\psi_{\flat}=\mathsf{p}_{M}\circ\mathsf{p}_{\mathfrak{g}[1]}$ to a dg-map by Remark A.1. In addition, we have to construct a dg-map
@@ 4.3 Construction of adjustment groupoids
To construct this map, we note that there is a pullback diagram of graded manifolds
@@
where $\mathsf{p}_{1,2}$ are the projections on the first and second summand in the Whitney sum bundle $\mathsf{t}^{*}\mathfrak{g}[1]\oplus\mathrm{T}[1]\mathsf{G}$, and $\mathsf{p}_{\mathsf{t}}\colon\mathsf{t}^{*}\mathfrak{g}[1]\rightarrow\mathfrak{g}[1]$ is the evident projection. We now define a map [^37] See Section A for further details on the notation.
@@
$\left(\prec^{\sharp}(\xi)\right)(\hat{\nu};g,\gamma)\coloneqq(\xi\circ\mathsf{p}_{\mathsf{s}})\left(\mathrm{Ad}\left(g^{-1},\hat{\nu}\right)+\vartheta^{\rm tot}(g,\gamma)\right)$ (4.21)
@@
In particular, the map Weierstrass-$\wp$ sends $(x,\bar{x},\alpha,\bar{\alpha};g,\gamma)$ to $(x,\alpha;g,\gamma)$, where $(x,\bar{x})$ belongs to $\mathrm{T}[1]M$, $(x,\alpha)$ belongs to $\mathfrak{g}[1]$, and $(x,\bar{\alpha})$ belongs to $\mathfrak{g}[2]$.
Definition 4.10 (Right action in adjustment groupoid).
[…]
@@
by Remark A.1.
[…]
Lemma 4.11.
[…]
@@
agree. Note that $(\psi\circ\mathbin{\vartriangleleft})_{\flat}$ and $(\mathrm{d}\mathsf{s})_{\flat}$ are fully defined by $|(\psi\circ\mathbin{\vartriangleleft})_{\flat}|=\mathsf{s}:\mathsf{G}\rightarrow M$ and $|(\mathrm{d}\mathsf{s})_{\flat}|=\mathsf{s}:\mathsf{G}\rightarrow M$, and hence they agree.
@@
It remains to show that the third relation in (2.44b) holds, i.e. that consecutive actions amount to groupoid composition. This is implied by the fact that this identity holds for the map $\prec^{\sharp}$ due to
$\mathrm{Ad}\left(g^{-1}_{2},\mathrm{Ad}\left(g^{-1}_{1},\hat{\nu}\right)+\vartheta^{\rm tot}(g_{1},\gamma_{1})\right)+\vartheta^{\rm tot}(g_{2},\gamma_{2})$ (4.26)
$\hskip 113.81102pt=\mathrm{Ad}\left((g_{1}\circ g_{2})^{-1},\hat{\nu}\right)+\vartheta^{\rm tot}\bigl{(}(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})\bigr{)}$
for all $(g_{1},\gamma_{1})\in\mathrm{T}[1]\mathsf{G}$ and $(g_{2},\gamma_{2})\in\mathrm{T}[1]\mathsf{G}$ with $(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})$ defined and $\hat{\nu}\in\mathfrak{g}_{\mathsf{t}(g_{1})}$, which is true for Cartan connections $\mathcal{H}$; that is, $\vartheta^{\rm tot}$ is a multiplicative 1-form w.r.t.  $\rm Ad$ whose horizontal distribution is complementary to $\mathsf{s}^{*}\mathfrak{g}$, which implies
$\mathrm{Ad}\left(g^{-1}_{2},\vartheta^{\rm tot}(g_{1},\gamma_{1})\right)+\vartheta^{\rm tot}(g_{2},\gamma_{2})=\vartheta^{\rm tot}\bigl{(}(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})\bigr{)}~{},$ (4.27)
@@
Definition 4.12 (Pre-adjustment groupoid $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$).
The pre-adjustment groupoid $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$ is the dg-Lie groupoid
@@
for all $V\in\mathrm{T}[1]\mathfrak{g}[1]$ and $\Gamma_{1,2}\in\mathrm{T}[1]\mathsf{G}$ with $\psi(V)=\mathrm{d}\mathsf{t}(\Gamma_{1})$ and such that all compositions and actions are defined. The dg-map $\psi\colon\mathrm{T}[1]\mathfrak{g}[1]\rightarrow\mathrm{T}[1]M$ is the unique map arising from the lift of $\psi_{\flat}=\mathsf{p}_{M}\circ\mathsf{p}_{\mathfrak{g}[1]}:\mathrm{T}[1]\mathfrak{g}[1]\rightarrow M$ to a dg-map by Remark A.1
We are then interested in the following statement.
Lemma 4.13.
The dg-Lie groupoid $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$ is a pre-adjustment groupoid in the sense of Definition 4.7.
@@
Functoriality of $\Psi$ amounts to showing that $\Psi_{0}$ is equivariant with respect to the $\mathrm{T}[1]\mathsf{G}$ actions. Using again Remark A.1, we have to show that
@@
$\mathsf{p}_{\mathsf{s}}\left(\mathrm{Ad}\left(g^{-1},\hat{\nu}\right)+\vartheta^{\rm tot}(g,\gamma)\right)~{},$ (4.33)

### v3

@@ 3.3 Adjusted local Lie-algebroid-valued connections
$\left(\mathrm{d}^{\nabla}\zeta\right)^{\alpha}_{abc}-\zeta^{\alpha}_{[a|d}\mathtt{r}^{d}_{\beta}\zeta^{\beta}_{|bc]}=0~{}.$ (3.24c)
@@
See the introduction for an extended history of these conditions; as major references there are [KS15b, general summary of adjusted algebroid theories; but curvature condition phrased differently, not as exactness condition] and [Fis21b, the exactness condition of the curvature first appeared here].
@@
Remark 3.13 (Integrated adjustment).
From an integrated point of view and as we discussed earlier, the vanishing of the basic curvature gives rise to the notion of Cartan connections $\mathcal{H}$ on the Lie groupoid $\mathscr{G}$ whose connection 1-form we denote by $\vartheta^{\mathrm{tot}}$; as pointed out in [Fis22] the curvature condition integrates to [^32] For the sake of readability we simplified the notation regarding the base points.
$\left.\left(\mathrm{d}^{\mathsf{s}^{*}\nabla}\vartheta^{\mathrm{tot}}-\frac{1}{2}\left(\mathsf{s}^{*}t_{\nabla_{\rho}}\right)\left(\vartheta^{\mathrm{tot}}\stackrel{{\scriptstyle\wedge}}{{,}}\vartheta^{\mathrm{tot}}\right)\right)\right|_{g}=\mathrm{Ad}_{\mathcal{H}}\left(g^{-1},\left.\mathsf{t}^{!}\zeta\right|_{g}\right)-\left.\mathsf{s}^{!}\zeta\right|_{g}$ (3.25)
for all $g\in\mathscr{G}$, where the exclamation marks denote the pullback of forms, $t_{\nabla_{\rho}}$ is the torsion of $\nabla_{\rho}$, and
$\left(\left(\mathsf{s}^{*}t_{\nabla_{\rho}}\right)\left(\vartheta^{\mathrm{tot}}\stackrel{{\scriptstyle\wedge}}{{,}}\vartheta^{\mathrm{tot}}\right)\right)(V,W)\coloneqq 2\left(\mathsf{s}^{*}t_{\nabla_{\rho}}\right)\left(\vartheta^{\mathrm{tot}}(V),\vartheta^{\mathrm{tot}}(W)\right)$ (3.26)
for all $V,W\in\mathfrak{X}(\mathscr{G})$. It might be surprising to see the torsion in the curvature; this is due to covariantising the theory, see [KS15b, Fis21b]. In fact, in Eq. (3.2b) we also pointed out that the components of the torsion are appearing.
@@
for $\mathcal{N}\in\mathsf{NQMfd}$. Here, $\mathsf{p}_{\rm curv}$ is the projection (2.51), and $\mathrm{deg}(-)$ denotes the degree of the enclosed morphism.
Remark 3.16.
The space $\mathfrak{brst}^{\rm adj}(U_{X},\mathfrak{g}[1])$ defined above exists as a dg-manifold if the basic curvature $R_{\nabla}^{\rm bas}$ vanishes. This space is generated by (see Section 3.2) the coordinate functions on field space $(\underline{\phi},\underline{A}_{\mathfrak{g}})$ in degree zero, the coordinate functions on gauge parameter space $(\underline{c}_{\mathfrak{g}},\underline{c}_{M},\underline{\lambda})$ in degree one and the coordinate functions on higher gauge parameter space $\underline{\chi}$ in degree two, quotient by the ideal generated by the equations
$\uline{c}^{a}_{M}=0~{},~{}~{}\uline{\lambda}^{\alpha}=0~{},~{}~{}\uline{\chi}^{\alpha}=0~{},~{}~{}\tfrac{1}{2}\left(R_{\nabla}^{\rm bas}\right)^{\alpha}_{\beta\gamma a}\uline{c}_{\mathfrak{g}}^{\beta}\uline{c}_{\mathfrak{g}}^{\gamma}\uline{A}_{M}^{a}=0~{}.$ (3.29)
The first three equations are linear in generators, and the quotient is straight-forward. The last one, however, is not linear and the quotient by the ideal it generates does not lead to a free algebra. In other words, the adjusted BRST complex does not form a dg-manifold [^33] Here we ignore subtleties related to working with infinite dimensions..
@@
00 0 $=\mathrm{d}E^{a}+\mathtt{r}^{a}_{\alpha}\left(F^{\alpha}-\omega_{b\beta}^{\alpha}E^{b}A^{\beta}-\tfrac{1}{2}\zeta^{\alpha}_{bc}E^{b}E^{c}\right)-\frac{\partial\mathtt{r}^{a}_{\alpha}}{\partial m^{b}}A^{\alpha}E^{b}~{},$ (3.32)
@@ 4.1 Inner action groupoids of dg-Lie groupoids
The inner action groupoid $\mathscr{A}(\mathscr{G})$ of a dg-Lie groupoid $\mathscr{G}$ is the internal homomorphism groupoid [^35] Recall $\sigma_{\Theta}(*)$ from Definition 2.14. $\underline{\mathsf{dgGrpd}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ with objects $\underline{\mathsf{hom}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ and morphisms the natural isomorphisms between the functors corresponding to $\underline{\mathsf{hom}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ in $\mathsf{dgMfd}$.
@@
$\begin{gathered}F_{1}(x,\theta_{1},\theta_{3})=F_{1}(x,\theta_{1},\theta_{2})F_{1}(x,\theta_{2},\theta_{3})~{},\\ \mathsf{t}(F_{1}(x,\theta_{1},\theta_{2}))=F_{0}(x,\theta_{1})~{},~{}~{}~{}\mbox{and}~{}~{}~{}\mathsf{s}(F_{1}(x,\theta_{1},\theta_{2}))=F_{0}(x,\theta_{2})~{},\end{gathered}$ (4.1b)
and $F_{1}$ is hence of the form
@@
$\begin{gathered}\mathsf{t}(m,\nu;g,\gamma)=(m,\nu)~{},~{}~{}~{}\mathsf{s}(m,\nu;g,\gamma)=(m,\nu)\mathbin{\blacktriangleleft}(g,\gamma)~{},\\ (m,\nu;g_{1},\gamma_{1})\circ\bigl{(}(m,\nu)\mathbin{\blacktriangleleft}(g_{1},\gamma_{1});g_{2},\gamma_{2}\bigr{)}=\bigl{(}m,\nu;(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})\bigr{)}~{},\\ \mathsf{e}(m,\nu)=(m,\nu;\mathsf{e}_{m},0)~{},~{}~{}~{}(m,\nu;g,\gamma)^{-1}=\left((m,\nu)\mathbin{\blacktriangleleft}(g,\gamma);(g,\gamma)^{-1}\right)\end{gathered}$ (4.9b)
@@
$(m,\nu)\mathbin{\blacktriangleleft}(g,\gamma)=\mathsf{p}_{s}\Big{(}\mathrm{Ad}_{\mathcal{H}}\left(g^{-1};m,\nu\right)+\vartheta^{\rm tot}(g,\gamma)\Big{)}~{},$ (4.12)
where $\mathrm{Ad}_{\mathcal{H}}$ is the adjoint action of Definition 2.9, and $\mathsf{p}_{s}$ is the projection $\mathsf{s}^{*}\mathfrak{g}\rightarrow\mathfrak{g}$. Note that both summands depend on a choice of connection, but this dependence drops out in the sum.
@@ 4.2 Adjustment groupoids of Lie groupoids
Definition 4.7 (Adjustment groupoid).
Consider a Lie groupoid $\mathscr{G}$ with Lie algebroid $\mathfrak{g}[1]=\mathsf{Lie}(\mathscr{G})[1]$. Then an adjustment groupoid of $\mathscr{G}$ is a dg-Lie groupoid $\mathscr{A}^{\mathsf{W}}=(\mathscr{A}^{\mathsf{W}}_{1}\rightrightarrows\mathrm{T}[1]\mathfrak{g}[1])$ together with a functor $\Psi:\mathscr{A}(\mathscr{G})\rightarrow\mathscr{A}^{\mathsf{W}}$ such that we have the following pullback diagram [^37] Here and in the following, $(\mathrm{id},0)$ denotes the embedding of a (graded) manifold into its (grade-shifted) tangent bundle as zero section.:
@@
Note that in the above definition, we could have equivalently demanded a pullback diagram involving the source maps. In the description of principal $\mathscr{G}$-connections, the functor $\Psi$ will ensure that flat connections are correctly included in our description of general connections, and the pullback diagram will ensure that we have the correct amount of gauge freedom.
@@
4.3 Adjustment groupoids and Cartan connections
In this section, we show that adjustments of Lie groupoids correspond precisely to Cartan connections.
Theorem 4.8.
Adjustment groupoids of a Lie groupoid $\mathscr{G}$ (Definition 4.7), are classified by Cartan connections on $\mathscr{G}$ (Definition 2.6).
Proof.
Note that any weak equivalence of adjustment groupoids (via bibundles) is given by strict isomorphisms, i.e. functors. Because of the choice $\mathscr{A}^{\mathsf{W}}_{0}=\mathrm{T}[1]\mathfrak{g}[1]$, for any weak equivalence of adjustment groupoids $\phi\colon\mathscr{A}^{\mathsf{W}}\cong\mathscr{A}^{\mathsf{W}\prime}$, there is a weakly commutative diagram
[…]
By [BX11, Lem. 2.29], $\phi$ can be represented as a functor $\phi$ which is identity on the space of objects $\mathrm{T}[1]\mathfrak{g}[1]$. This allows us to split the proof into two arguments presented further below.
Given a Cartan connection, there is a right action of $\mathrm{T}[1]\mathsf{G}$ on $\mathrm{T}[1]\mathfrak{g}[1]$ in the sense of Definition 2.24, as we will show in Lemma 4.9, and the corresponding action groupoid $\mathscr{A}_{\mathcal{H}}^{\mathsf{W}}$ in the sense of Definition 2.25, is an adjustment groupoid, as we will show in Lemma 4.10.
Conversely, given an adjustment groupoid, we will construct a Cartan connection in Proposition 4.13. From the constructions, it is clear that the adjustment defined by the Cartan connection on $\mathscr{G}$ derived from an adjustment groupoid $\mathscr{A}^{\mathsf{W}}$ yields again $\mathscr{A}_{\mathcal{H}}^{\mathsf{W}}=\mathscr{A}^{\mathsf{W}}$. ∎
We start with the construction of the action mentioned in the above proof. In the following, let $\mathscr{G}=(\mathsf{G}\rightrightarrows M)$ be a general Lie groupoid and $\mathfrak{g}[1]$ its Lie algebroid, regarded as a grade-shifted vector bundle. Furthermore, let $\mathcal{H}$ be a Cartan connection on $\mathscr{G}$.
The dg-right action of $\mathrm{T}[1]\mathscr{G}$ on $\mathrm{T}[1]\mathfrak{g}[1]$ is defined along a dg-map $\psi\colon\mathrm{T}[1]\mathfrak{g}[1]\rightarrow\mathrm{T}[1]M$, and the relevant map here is the unique map arising from the lift of [^38] As before, $\mathsf{p}_{M}$ denotes a projection onto the base $M$ of a fiber bundle $E\rightarrow M$. $\psi_{\flat}=\mathsf{p}_{M}\circ\mathsf{p}_{\mathfrak{g}[1]}$ to a dg-map by Remark A.1. In addition, we have to construct a dg-map
@@ 4.3 Adjustment groupoids and Cartan connections
We now note that there is a pullback diagram of graded manifolds
@@
where $\mathsf{p}_{1,2}$ are the projections on the first and second summand in the Whitney sum bundle $\mathsf{t}^{*}\mathfrak{g}[1]\oplus\mathrm{T}[1]\mathsf{G}$, and $\mathsf{p}_{\mathsf{t}}\colon\mathsf{t}^{*}\mathfrak{g}[1]\rightarrow\mathfrak{g}[1]$ is the evident projection. We can then define a map [^39] See Section A for further details on the notation.
@@
$\left(\prec^{\sharp}(\xi)\right)(\hat{\nu};g,\gamma)\coloneqq(\xi\circ\mathsf{p}_{\mathsf{s}})\left(\mathrm{Ad}_{\mathcal{H}}\left(g^{-1},\hat{\nu}\right)+\vartheta^{\rm tot}(g,\gamma)\right)$ (4.21)
@@
In particular, the map Weierstrass-$\wp$ sends $(x,\bar{x},\alpha,\bar{\alpha};g,\gamma)$ to $(x,\alpha;g,\gamma)$, where $(x,\bar{x})$ belongs to $\mathrm{T}[1]M$, $(x,\alpha)$ belongs to $\mathfrak{g}[1]$, and $(x,\bar{\alpha})$ belongs to $\mathfrak{g}[2]$. This gives us what we need:
Lemma 4.9.
Consider the map
@@
defined as the unique lift of $\mathbin{\vartriangleleft}_{\flat}$ from (4.23) by Remark A.1. This map defines a right action of $\mathrm{T}[1]\mathscr{G}$ on $\mathrm{T}[1]\mathfrak{g}[1]$.
@@
agree. Note that both $(\psi\circ\mathbin{\vartriangleleft})_{\flat}$ and $(\mathrm{d}\mathsf{s})_{\flat}$ are fully defined by $|(\psi\circ\mathbin{\vartriangleleft})_{\flat}|=\mathsf{s}:\mathsf{G}\rightarrow M$ and $|(\mathrm{d}\mathsf{s})_{\flat}|=\mathsf{s}:\mathsf{G}\rightarrow M$, and hence they agree.
@@
The third relation in (2.45b), i.e. consecutive actions amount to groupoid composition, is implied by the fact that this relation holds for the map $\prec^{\sharp}$. This is seen from
$\mathrm{Ad}_{\mathcal{H}}\left(g^{-1}_{2},\mathrm{Ad}_{\mathcal{H}}\left(g^{-1}_{1},\hat{\nu}\right)+\vartheta^{\rm tot}(g_{1},\gamma_{1})\right)+\vartheta^{\rm tot}(g_{2},\gamma_{2})$ (4.26)
$\hskip 113.81102pt=\mathrm{Ad}_{\mathcal{H}}\left((g_{1}\circ g_{2})^{-1},\hat{\nu}\right)+\vartheta^{\rm tot}\bigl{(}(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})\bigr{)}$
for all $(g_{1},\gamma_{1})\in\mathrm{T}[1]\mathsf{G}$ and $(g_{2},\gamma_{2})\in\mathrm{T}[1]\mathsf{G}$ with $(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})$ defined and $\hat{\nu}\in\mathfrak{g}_{\mathsf{t}(g_{1})}$, which is true for Cartan connections $\mathcal{H}$. That is, $\vartheta^{\rm tot}$ is a multiplicative 1-form with respect to $\rm Ad_{\mathcal{H}}$, whose horizontal distribution is complementary to $\mathsf{s}^{*}\mathfrak{g}$, which implies
$\mathrm{Ad}_{\mathcal{H}}\left(g^{-1}_{2},\vartheta^{\rm tot}(g_{1},\gamma_{1})\right)+\vartheta^{\rm tot}(g_{2},\gamma_{2})=\vartheta^{\rm tot}\bigl{(}(g_{1},\gamma_{1})\circ(g_{2},\gamma_{2})\bigr{)}~{},$ (4.27)
@@
Lemma 4.10.
Consider the action dg-groupoid $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$ given by
@@
for all $V\in\mathrm{T}[1]\mathfrak{g}[1]$ and $\Gamma_{1,2}\in\mathrm{T}[1]\mathsf{G}$ with $\psi(V)=\mathrm{d}\mathsf{t}(\Gamma_{1})$ and such that all compositions and actions are defined. Moreover, the dg-map $\psi\colon\mathrm{T}[1]\mathfrak{g}[1]\rightarrow\mathrm{T}[1]M$ is the unique map arising from the lift of $\psi_{\flat}=\mathsf{p}_{M}\circ\mathsf{p}_{\mathfrak{g}[1]}:\mathrm{T}[1]\mathfrak{g}[1]\rightarrow M$ to a dg-map by Remark A.1
This dg-Lie groupoid is an adjustment groupoid in the sense of Definition 4.7.
@@
[…]
@@
$\mathsf{p}_{\mathsf{s}}\left(\mathrm{Ad}_{\mathcal{H}}\left(g^{-1},\hat{\nu}\right)+\vartheta^{\rm tot}(g,\gamma)\right)~{},$ (4.33)